OpenStudy (anonymous):

The solution set of the equation |x-14| = |x +14| is?

OpenStudy (shubhamsrg):

0 is the only soln.. x-14 = x+14 and x-14 = (-x+14)

OpenStudy (anonymous):

Isn't when we solve a mod we do like this: |x-14| = |x +14| x-14 = \[\pm \] (x +14) So,

OpenStudy (shubhamsrg):

i did nothing different mam..

OpenStudy (anonymous):

um.............u l have three cases if u wanna do it that way

OpenStudy (anonymous):

x-14 = x+14 Then expand with negative X-14 = -x-14 ??? @shubhamsrg

OpenStudy (shubhamsrg):

i meant that you'll have to solve both eqns separately.. both will yield x=0

OpenStudy (anonymous):

I dont get it :(

OpenStudy (shubhamsrg):

hmmn.. you have this,, firstly,,x+14 = x-14 solving gives no value of x..(am sorry,,not 0) next eqn is (x+14)= -(x-14) now you'll x=0 was it clearer this time mam ?

OpenStudy (anonymous):

Yep Thank you so much, but why are you calling me mam? Please dont do :)

OpenStudy (shubhamsrg):

lol.. XD sorry MAM! :P

OpenStudy (anonymous):

another way you might consider is looking at the graphs of the left side and right side.... they only intersect at x= 0....